Julie and Eric row their boat (at a constant speed) 32 miles downstream for 4 hours, helped by the current. Rowing at the same rate, the trip back against the current takes 8 hours. Find the rate of the current.
Answer by TimothyLamb(4379)

s = d / t
d = s * t
---
b + w = 32/4 = 8 mph
b - w = 32/8 = 4 mph
b - w = 4 mph
b = 4 + w
---
b + w = 8
(4 + w) + w = 8
4 + 2w = 8
---
w = 2 mph
b = 6 mph
---
Answer:
the water current is moving at 4 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php